A stone is dropped from the top of a tower and travels 24.5 in the last second of its journey. The height of the tower

Grade:12th pass

1 Answers

Vikas TU
7 years ago
From the eqn. 
s1  = 0 + 0.5gt^2.............(1)
s2 = 0 + 0.5g(t-1)^2........(2)
s1 – s2 = 5t^2 – 5(t-1)^2 = 24.5
on solving we get,
t = 2.95 seconds
and
s1 = 43.5 meter.